OpenStudy (anonymous):

please help!!! Write the equation of the line that is perpendicular to y= 1/4x and passes through the point (2,-5)

OpenStudy (anonymous):

How do you find the slope that is perpendicular?

OpenStudy (mathmate):

slopes s and s' are perpendicular if s*s' = -1

OpenStudy (anonymous):

not sure @gypsy1274

OpenStudy (anonymous):

If you draw each line on a graph, it will make an X. So, if the original slope is positive, what will be the sign of the perpendicular slope?

OpenStudy (anonymous):

negative

OpenStudy (anonymous):

Ok. So, You know that the slope will be negative. Will it still be 1/4? Draw a quick sketch.

OpenStudy (anonymous):

wouldn't it be -1/4

OpenStudy (anonymous):

No. Poor drawing skills, but it gives you a general idea. 

hero (hero):

@xoxox123

OpenStudy (anonymous):

@Hero its wrong ??

hero (hero):

y= 1/4x and passes through the point (2,-5) If the line is perpendicular to the given line, then it will have a slope of -4 so y = -4x + b Now insert the point to find b: -5 = -4(2) + b -5 = -8 + b 8 - 5 = b 3 = b So the new formula is y = -4x + 3

hero (hero):

Sorry, now it is correct.

OpenStudy (anonymous):

no.. but 2 other people told me that answer too.

OpenStudy (anonymous):

@Hero

hero (hero):

Trust me, it is correct.

OpenStudy (anonymous):

Hero's second answer is correct.

hero (hero):

I only have one answer lol

OpenStudy (mathmate):

Unless there is a typo in your question!

hero (hero):

^What @mathmate said

OpenStudy (mathmate):

Do you mean y=(1/4)*x

OpenStudy (texaschic101):

I told you @xoxox123 ...we were not wrong

OpenStudy (mathmate):

If the question is y=(1/4)x, and the point through which the perpendicular passes, then Hero is right.

OpenStudy (mathmate):

(2,-5) is the point, I mean.

OpenStudy (texaschic101):

Its got to be the website xoxo123....you are getting the same answer from everyone

OpenStudy (texaschic101):

good idea :)